1. Warmer air moved into the city last night, raising the temperature from a low of −7° overnight to a high of 19° this morning. What was the total increase in temperature?''.
Answer by Boreal(15235)

The increase was 7 degrees to 0 and another 19 degrees after that, for 26 degrees.
19-(-7)=19+7=26 deg